OpenStudy (anonymous):

How can I find all the value of sin(3θ)=(1/2) in the interval [0, 2pi]

terenzreignz (terenzreignz):

Well, first, what angles do you know have a sine of 1/2?

OpenStudy (anonymous):

pi/6 and (2pi)/3

terenzreignz (terenzreignz):

\[\large \frac{2\pi }{3}\]? I don't think so :P

OpenStudy (anonymous):

Blah, I meant to say (5pi)/6

OpenStudy (anonymous):

I know that there is supposed to be 6 values of θ, I just don't know how to get to that point.

terenzreignz (terenzreignz):

Okay, then think of six angles whose sines are all 1/2. Crucial thing is not to confine yourself in \(\large [0,2\pi]\) but rather, triple that... into \[\Large [0,6\pi]\] and see what we get

OpenStudy (anonymous):

OH! pi/6, (5pi)/6, (13pi)/6, (17pi)/6, (25pi)/6, (29pi)/6

OpenStudy (anonymous):

I then take all these divide by three

OpenStudy (anonymous):

So my final answers would be pi/18, (5pi)/18, (13pi)/18, (17pi)/18, (25pi)/18, (29pi)/18

OpenStudy (anonymous):

Are those correct?